. A current of 2.0 A flows in a circuit. How much charge passes a point in the circuit in TOS
1
Expert's answer
2020-06-17T09:22:58-0400

Current in the circuit:

"I=\frac{\Delta q}{\Delta t};"

Charge per unit time:

"q=\Delta q=I\cdot t;"

Charge per one second:

"q=2A\cdot 1s=2 coulomb;"

Number of electrons per second:

"N=\frac{q}{e}=\frac{2}{1.6\cdot 10^{-19}}=125\cdot 10^{17};"